OpenStudy (anonymous):

For what values of x does the series 1+2^x+3^x+4^x+...+n^x+... converge? (Answer: x<-1)

OpenStudy (anonymous):

my guess is \(x<1\)

OpenStudy (anonymous):

But the answer says x<-1, are you sure your answer is right? Can you show the work?

OpenStudy (anonymous):

oh my guess was wrong, sorry ignore that one

OpenStudy (kc_kennylau):

\[\large\begin{array}{rcl} \left|\lim_{n\rightarrow\infty}\frac{(n+1)^x}{n^x}\right|&<&1\\ \left(\lim_{n\rightarrow\infty}\frac{(n+1)^x}{n^x}\right)^2&<&1\\ \lim_{n\rightarrow\infty}\frac{(n+1)^{2x}}{n^{2x}}&<&1\\ \lim_{n\rightarrow\infty}[\ln(n+1)^{2x}-\ln n^{2x}]&<&0\\ \lim_{n\rightarrow\infty}[2x\ln(n+1)-2x\ln n]&<&0\\ \lim_{n\rightarrow\infty}2x[\ln(n+1)-\ln n]&<&0\\ \end{array}\]Wait this seems be not the approach correct...

OpenStudy (kc_kennylau):

Forgive me English of mine

OpenStudy (kc_kennylau):

@satellite73 you think something?

OpenStudy (kc_kennylau):

@Loser66 and about you?

OpenStudy (anonymous):

i can't seem to figure out how to get the limit less than 1

OpenStudy (kc_kennylau):

Maybe the test wrong I using?

OpenStudy (anonymous):

i think the log is the way to go

OpenStudy (kc_kennylau):

@Idealist What think-you?

OpenStudy (anonymous):

\[(\frac{n+1}{n})^x<1\] \[x\ln(n+1)-\ln(x)<0\] but i keep going round in circles

OpenStudy (kc_kennylau):

But same is that of mine?

OpenStudy (anonymous):

\[x\ln(\frac{n+1}{n})<0\] yes same

OpenStudy (kc_kennylau):

And limit approach zero?

OpenStudy (kc_kennylau):

Gives you 0x<0?

OpenStudy (kc_kennylau):

That not makes sense...

OpenStudy (anonymous):

having a brain fart i think the key is solving \[x\ln(\frac{n+1}{n})<0\] for \(x\)

OpenStudy (kc_kennylau):

But that gives x*0<0 which not makes sense...

OpenStudy (anonymous):

which wolfram tells me the solution is \(x<-1\)

OpenStudy (kc_kennylau):

interesting

OpenStudy (anonymous):

But what did wolfram do?

OpenStudy (anonymous):

dunno i must be screwing up somewhere

OpenStudy (anonymous):

What do you mean by false?

OpenStudy (anonymous):

if \(x=-1\) you get \(\sum\frac{1}{n}\) which is well known to diverge

OpenStudy (kc_kennylau):

google "harmonic series" for more info :)

OpenStudy (anonymous):

if \(x<-1\) say \(x=-(1+\epsilon)\) you get \[\sum\frac{1}{n^{1+\epsilon}}\] where \(\epsilon>1\) and this converges, probably easiest to use the integral test

OpenStudy (anonymous):

i meant where \(\epsilon>0\)

OpenStudy (kc_kennylau):

But how do we know that it converges if x>-1?

OpenStudy (kc_kennylau):

diverges*

OpenStudy (anonymous):

if it diverges at \(x=-1\) it certainly diverges for \(x>-1\)

OpenStudy (anonymous):

in any case if \(x=-1+\epsilon\) where \(\epsilon>0\) then you get \[\sum\frac{1}{n^{1-\epsilon}}\] and that diverges again by the integral test

OpenStudy (anonymous):

got off on the wrong foot by thinking of ratio test, rather than harmonic series etc

OpenStudy (anonymous):

like i said, brain fart

OpenStudy (anonymous):

HAHAHA!